Jacob can do 4 jobs in 5 hours. When Kristian helps, together they manage to do 56 jobs in 20 hours. How long would it take Kristian to do 1 job?
Answer by VFBundy(438)

Rate per hour it takes Jacob to do one job = 4/5

Rate per hour it takes Jacob and Kristian to do one job = 56/20 = 14/5


Jacob's rate of jobs per hour + Kristian's rate of jobs per hour = 14/5

4/5 + Kristian's rate of jobs per hour = 14/5

Kristian's rate of jobs per hour = 14/5 - 4/5

Kristian's rate of jobs per hour = 14/5 - 4/5 = 10/5 = 2

So, we now know that Kristian can do 2 jobs per hour.

If Kristian can do 2 jobs per hour, this means it will take him 0.5 hours to do one job.